hxkdkyc:

the student council is having a bake sale. they spend $40 for the ingredients to make brownies and $1.50 to make a dozen cookies. the brownies sell for $6 per dozen and the cookies sell for $7.50 per dozen. the student council sells a total of $195 in bake goods

doggofroyo:

Where's the question?

hxkdkyc:

e write a linear equation with two equations and two variables represent the situation

doggofroyo:

Oh.

mikewwe13:

Let's first define some variables to help us solve the problem: - Let B be the number of dozens of brownies sold. - Let C be the number of dozens of cookies sold. From the problem statement, we know that: - The cost of the ingredients for the brownies is $40. - The cost of making a dozen cookies is $1.50. - The selling price of a dozen brownies is $6. - The selling price of a dozen cookies is $7.50. - The total revenue from the bake sale is $195. Using these variables and equations, we can set up a system of two equations to solve for the unknowns B and C: Equation 1: B*6 + C*7.5 = 195 This equation represents the total revenue from the sale of brownies and cookies. Equation 2: B*40/12 + C*1.5 = total cost This equation represents the total cost of making the brownies and cookies. To solve for B and C, we need to first find the total cost of making the brownies and cookies. Total cost = cost of brownies + cost of cookies Total cost = $40 + C*$1.50 Total cost = $40 + $1.5C Substituting this expression for the total cost into Equation 2, we get: B*40/12 + C*1.5 = $40 + $1.5C Simplifying this equation, we get: 10B + 3C = 80 Now we have two equations with two unknowns: 10B + 3C = 80 6B + 7.5C = 195 We can solve this system of equations using any method of our choice (such as substitution or elimination). One possible method is to multiply the first equation by 2 and subtract it from the second equation: 6B + 7.5C = 195 - (20B + 6C = 160)


-14B + 1.5C = 35 Solving for C in this equation, we get: 1.5C = 14B + 35 C = (14/1.5)B + 35/1.5 C = 9.33B + 23.33 Substituting this expression for C into one of the earlier equations, we get: 10B + 3(9.33B + 23.33) = 80 10B + 28B + 70 = 80 38B = 10 B = 10/38 B ≈ 0.263 So the student council sold approximately 0.263 dozen brownies. Using this value of B, we can solve for C using the expression we derived earlier: C = 9.33B + 23.33 C ≈ 9.33(0.263) + 23.33 C ≈ 25.49 So the student council sold approximately 25.49 dozen cookies. Therefore, the student council sold approximately 0.263 dozen brownies and 25.49 dozen cookies at their bake sale.